Multiply the following using a diagrammatical representation:
1. 2/3 *6/8
2. 3/4 * 4/5
3. 3/5 * 2/3
4. 3/4 * 2/2
1
Expert's answer
2020-10-21T15:06:42-0400


1. The boxes coloured black represents "\frac{2}{3}" and the boxes coloured red represents "\frac{6}{8}" .


N.B: Some boxes contain both colours and some do not contain any colour.


Number of boxes that contain both colours = 12

Total number of boxes = 24

"\frac{2}{3} × \frac{6}{8}" = "\frac{12}{24} = \frac{1}{2}"





2. The boxes coloured blue represents "\frac{3}{4}" and the boxes coloured red represents "\frac{4}{5}" .


N.B: Some boxes contain both colours and one box does not contain any colour.


Number of boxes that contain both colours = 12

Total number of boxes = 20

"\frac{3}{4} × \frac{4}{5} = \frac{12}{20} = \frac{3}{5}"





3. The boxes coloured blue represents "\frac{2}{3}" and the boxes coloured red represents "\frac{3}{5}" .


N.B: Some boxes contain both colours and some do not contain any colour.


Number of boxes that contain both colours = 6

Total number of boxes = 15

"\frac{3}{5} × \frac{2}{3} = \frac{6}{15} = \frac{2}{5}"





4. The boxes coloured blue represents "\frac{2}{2}" and the boxes coloured red represents "\frac{3}{4}" .


N.B: Some boxes contain both colours.


Number of boxes that contain both colours = 6

Total number of boxes = 8

"\frac{3}{4} × \frac{2}{2} = \frac{6}{8} = \frac{3}{4}"